OpenStudy (hayleymeyer):

help pweezz!! What is the difference in volume between a sphere with a radius of 12 centimeters and a sphere with a diameter of 12 centimeters? a. 6,330.24 cm3 c. 3,617.28 cm3 b. 5,994.36 cm3 d. 904.32 cm3

OpenStudy (hayleymeyer):

I thought B.

OpenStudy (anonymous):

the volume is 7,238.2 cubic cm

OpenStudy (anonymous):

for the radius of 12 sphere

OpenStudy (anonymous):

\[difference~ \in~ volume=\frac{ 4 }{ 3 }\pi \left( 12^3-6^3 \right)=\frac{ 4 }{ 3 }\pi *6^3\left( 2^3-1 \right)\] \[=\frac{ 4 }{ 3 }*216\left( 8-1 \right)\pi=288*7 \pi=?\] i can't guess, you can calculate.

OpenStudy (hayleymeyer):

thanks everyone

OpenStudy (anonymous):

B is not correct